OpenStudy (anonymous):

use a substitution to solve the system of equations. 3x-3y=30 , -2x+8y=-26

OpenStudy (anonymous):

answers are a. (4,-6) b. (9,-1) c. (17,1) d. infinitely many solutions

OpenStudy (imtant):

Firstly you need to multiply both eq. so that x or y have the same values.Then you can add or substract it. For example: both equation I multiply by 2.So the equation should be like this, 6x-6y=60 -6x+24y=-78

OpenStudy (imtant):

You can see that the value of x is same right? So you can add 'plus sign' to the first eq. 6x-6y=60 Then you will get 18y=-18 which y=-1

OpenStudy (anonymous):

ok what do i do next ?

OpenStudy (imtant):

So last step you need to substitute the value of y into the any the first eq. before you multiply by 2. For example: 3x-3y=30 So put the y value into the eq. You will get the x value. Ok?

OpenStudy (anonymous):

So x is 9 ?

OpenStudy (imtant):

Yes,that's it.

OpenStudy (anonymous):

Thanks

OpenStudy (imtant):

Wc.